OpenStudy (anonymous):

how to solve 3w^2 + 22w + 7 = 0 ?

OpenStudy (anonymous):

quadratic formula.

OpenStudy (anonymous):

I have no idea what that is yet... sorry.

OpenStudy (displayerror):

For \[ax^2 + bx + c\] The quadratic formula is given by \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\]

OpenStudy (anonymous):

yeah so a=3, b = 22, c = 7

OpenStudy (displayerror):

Okay, so plug it in!

OpenStudy (anonymous):

I have no idea how to even begin... My teacher had just thrown the work at us without teaching us or the quadratic formula..

OpenStudy (anonymous):

Just to show that we could try to solve it on our own... but really im so confused. :/

OpenStudy (anonymous):

\[\frac{ -22+\sqrt{(22^2)-4(3)(7)} }{ 2*3 }\]

OpenStudy (anonymous):

its just a plug and chug with the equation.

OpenStudy (anonymous):

okay thank you :) I'll see how it goes..

OpenStudy (anonymous):

Good luck!